OpenStudy (anonymous):

Prove The Identity: 1-tan^4 θ = 2 sec^2 θ - sec ^4 θ

OpenStudy (paxpolaris):

factor both sides:

OpenStudy (anonymous):

\[ 1- \tan^4 x =(1-\tan^2 x)(1+\tan^2 x)=(1-\tan^2 x) \sec^2 x=\\ \left( 1- \frac {\sin^2 x}{\cos^ x} \right)\sec^2 x \] The rest should be easy

OpenStudy (anonymous):

Given: 3x - 2 ≤ 2x + 1. Choose the solution set. {x | xR, x ≤ -3} {x | xR, x ≤ -1} {x | xR, x ≤ 1} {x | xR, x ≤ 3}

OpenStudy (anonymous):

i have no clue about this question, i'm only in 9th grade math. So, are the solutions the answers?

OpenStudy (anonymous):

\[1- \tan^4 x =(1-\tan^2 x)(1+\tan^2 x)=(1-\tan^2 x) \sec^2 x=\\ \left( 1- \frac {\sin^2 x}{\cos^2 x} \right)\sec^2 x=\\ \left( \frac{\cos^2 x- \sin^2 x}{\cos^2 x}\right)\sec^2 x=\\ \left(\frac{2\cos^2 x -1}{\cos^2 x}\right)\sec^2 x=\\ \left(2 -\sec^2 x\right)\sec^2 x=\\ 2 \sec^2 x - \sec^4 x \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@Camoguy937 did you understand my solution?

OpenStudy (anonymous):

yes, thanks!

OpenStudy (anonymous):

yw